Complete the table below for the missing amounts

A

B

C

Number of units

4,512

units

(d)

2,112

units

Sale price per unit

$245

$192

(g)

Variable costs per unit

(a)

248

3,014

Contribution margin per unit

$98

(e)

(h)

Total contribution margin

(b)

$(1,324,400)

(i)

Contribution margin ratio

(c)

(f)

45

%

Compute the missing​ information, starting with scenario​ A, then for scenarios B and C.

A

Number of units

4,512

units

Sale price per unit

$245

Variable costs per unit

?

Contribution margin per unit

$98

Total contribution margin

?

Contribution margin ratio

?

%
Answer #1

Complete the table below for the missing amounts

A

B

C

Number of units

4,512

units

(d)1324400/56 = 23650

2,112

units

Sale price per unit

$245

$192

(g)2466+3014 = 5480

Variable costs per unit

(a) (245-98) = 147

248

3,014

Contribution margin per unit

$98

(e) 192-248 = -56

(h)3014*45/55 = 2466

Total contribution margin

(b) 98*4512 = 442176

$(1,324,400)

(i)2112*2466 = 5208192

Contribution margin ratio

(c)98/245 = 40%

(f) 56/192 = -29%

45

%